A pendulum clock is 5 sec fast at temperature of 15 °C and 10 sec slow at a temperature of 30 °C. At what temperature does it give the correct time?

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
The change in time period of a pendulum and the temperature can be explained with
delta t = 1/2alpha Delta t 
therefore the equation will form as
 
5/10= T-15 / 30-T ,
from where we would get T=20deg at which we would get the correct time
Good Luck